OpenStudy (anonymous):

what is the slope ;line of 7y+5x=10

OpenStudy (anonymous):

bring the equation to the form: y=mx+n in this form m = slope .

OpenStudy (anonymous):

ok so tell me what is the slope?

OpenStudy (anonymous):

7y=-5x+10 y = (-5/7)x + 10/7 m=-5/7

OpenStudy (anonymous):

thank u